# Grounding Synthetic Data Evaluations of Language Models in Unsupervised Document Corpora

### v1

@@ 4.2 Question Generation Models
To explore the impact of model capability on successful generation of synthetic data benchmarks we tested a variety of language models during the synthetic benchmark generation phase.All open-weight models are available through HuggingFace [47] and the closed-weight models were accessed through paid APIs. Additional results are reported for an ensemble of the bolded models. During benchmark creation the same model is used both for topic extraction and question generation.
@@
    google/gemma-3-4b-it [48]
@@
    google/gemma-3-12b-it [48]
@@
    google/gemma-3-27b-it [48]
@@
    meta-llama/Llama-3.2-3B-Instruct [49]
  5. 5.
    meta-llama/Llama-3.1-8B-Instruct [50]
  6. 6.
    meta-llama/Llama-3.3-70B-Instruct [50]
  7. 7.
    meta-llama/Llama-4-Maverick-17B-128E-Instruct [51]
  8. 8.
    openai/o4-mini [52]
  9. 9.
    openai/gpt-4.1 [52]
  10. 10.
    openai/gpt-4.1-mini [52]
  11. 11.
    microsoft/phi-4 [53]
  12. 12.
    Qwen/Qwen2.5-7B-Instruct [54]
4.3 Evaluating Against Human Questions
The reformatted human annotated question/answer pairs are used to compare the grounded synthetic data benchmark performance against the human. Ideally, the synthetic benchmark from grounding documents would produce the same LM capability measurements as the human curated benchmark for any given model under evaluation. Thus, to validate our approach, various LMs of differing sizes are evaluated on both the human and synthetic benchmarks, and we expect the benchmark scores will correlate between original and novel dataset across all evaluated models. Comparison of pure benchmark evaluation accuracy is performed using the Pearson accuracy correlation between human and synthetic benchmark accuracy numbers. Evaluating the comparative rank ordering of various models under test uses the Spearman rank correlation of relative model performance on the benchmark. These two measures compare both absolute performance alignment and relative rank order preservation/stability.
4.4 Generated Question Quality Characterization
Per-question quality features and benchmark dataset-level features are computed to determine the overall quality of the synthetic benchmark. LMs are used to extract the following information about each generated question:
  1. 1.
    question token count: number of LM tokens in the question
  2. 2.
    question clarity: how clear and comprehensible the question is in isolation without the associated context
  3. 3.
    question groundedness: how grounded the question is in the context
  4. 4.
    answer correctness: how correct the answer is given the context
@@ 4.4 Generated Question Quality Characterization
Owing to the low level of human oversight and annotation time for creating the synthetic data benchmarking, LMs are leveraged to extract these somewhat imprecise features. This relies on the idea that with enough guidance and prompting LMs can be good alternatives to human evaluation [28, 29]. This is also similar to how standard practice LLM-as-a-Judge is used for grading open-ended benchmark responses [30, 31]. In addition to per-question measurements, the question diversity is measured over the whole benchmark as the average cosine similarity between all questions (using sentence-transformers/all-MiniLM-L6-v2 as the embedding model). The values of answer_correctness and answer_explanation_validity are used to determine which question generating models are incapable of producing diagnostically useful questions.
@@ 5 Results
5.1 Generated Question Quality
Understanding the quality of synthetically generated questions is useful to evaluate the applicability of this approach for system capability evaluation. Unless otherwise specified, this work uses Llama-4-Maverick-17B-128E-Instruct as the feature extractor from the text of questions and answers.
5.1.1 Answer Correctness
For a synthetic benchmark, it is crucial that generated answers are accurate and correct. Fortunately, the task of generating correct answers is simplified by the grounding context, which ideally contains the necessary information for the model to reformat into a question and correct answer. To verify generated answer accuracy, an LM examined each question-answer pair alongside the full context chunk and evaluated both the answer correctness and the validity of the generating model’s explanation for the answer (both on a scale of 1-10). The results, shown in fig. 3, demonstrate that most non-small models achieve an average correctness score greater than $9.5$ across all datasets. In contrast, smaller and less capable models showed a marked decline in generated answer correctness. Throughout our experimentation, we observed a strong correlation between model size/capability and desirable properties for synthetic data generation. However, we also found some notable exceptions (like gpt-4.1), which will be discussed later.
@@ 5.1.1 Answer Correctness
Figure 3: Synthetic data generated question correctness and explanation validity graded by Llama-4-Maverick-17B-128E-Instruct (potentially not the generating model) on a scale of 1 to 10.
5.1.2 Question Clarity and Groundedness
Question clarity is inherently tied to difficulty, as simple questions have the potential to be clearer and easier to answer. Conversely, a needlessly opaque question is harder to answer without providing additional diagnostic relevance, though a model’s ability to reason through and understand more opaque questions might indicate general capability. This study found that most generated questions have reasonable clarity without significant variation between the generating models. Most generated synthetic questions are longer and more complicated than the original questions. This increased average question complexity combined with similar levels of question clarity suggests most synthetic questions are well written.
How grounded in the source material (context chunk) a given question is indicates how much outside knowledge and information both the generation and evaluation models must leverage to correctly interpret what is being asked and the correct answer. A low grounding score indicates the question/answer tuple cannot be found within the context chunk, or that significant outside information is required. A high grounding score means all necessary information is found within the context chunk. This indicates both that the question is well aligned with the context chunk information and that the answer is available in that context. Therefore the generated question is testing for information contained in the context chunk. Figure 4 highlights the question clarity and groundedness scores for various question generating models. While most generating models have similar clarity values, the trends in groundedness score are more complex. Some capable models (like Llama-4-Maverick-17B-128E-Instruct) score well on groundedness, other capable models (like o4-mini) score lower, with smaller models in between (gemma-3-4b-it). We observed two failure modes with respect to grounding: Either the model hallucinates content absent from and not relevant to the context chunk (phi-4 commonly did this), or the generating model uses knowledge in its weights to expand upon the information presented in the context (o4-mini). For example, phi-4 generated the following question from the provided context:
2"context": "Victoria later described her childhood as \"rather melancholy\". Her mother was extremely protective, and Victoria was raised largely isolated from other children under the so-called \"Kensington System\", an elaborate set of rules and protocols devised by the Duchess and her ambitious and domineering comptroller, Sir John Conroy, who was rumoured to be the Duchesss lover. The system prevented the princess from meeting people whom her mother and Conroy deemed undesirable (including most of her fathers family), and was designed to render her weak and dependent upon them. The Duchess avoided the court because she was scandalised by the presence of King Williams bastard children, and perhaps prompted the emergence of Victorian morality by insisting that her daughter avoid any appearance of sexual impropriety. Victoria shared a bedroom with her mother every night, studied with private tutors to a regular timetable, and spent her play-hours with her dolls and her King Charles spaniel, Dash. Her lessons included French, German, Italian, and Latin, but she spoke only English at home."
Interestingly enough, models of middle size/capability perform well on groundedness, just reformatting the information presented to them in the context.
@@ 5.1.2 Question Clarity and Groundedness
Figure 4: Generated question clarity and groundedness compared to the original questions.
5.1.3 Impact of Question Length
The synthetic data questions in this study tend to be much longer than the original human-written questions. The generating model is encouraged to include all information required to disambiguate what the synthetic question is asking about, causing the questions to be longer and more detailed questions than what humans might write. Figure 5 (left) highlights this trend. Some generating LMs take this to an extreme, causing the synthetic questions to become easier for models to answer correctly. We hypothesize this happens because with more detail in the question, it provides the LM under evaluation more information to pattern match against text seen in the training phase, increasing the likelihood that the LM retrieves memorized patterns relevant to the question. This effect is seen in Figure 5 (right) where the LM benchmark evaluation accuracy over-performance (synthetic question accuracy - original accuracy) is plotted. There is a loose trend where the longer a question the more models over-perform the $y=0$ expected trend line.
@@ 5.1.3 Impact of Question Length
(a) Generated question token count by generating model. Most models are within $20\%$ of the human length.
@@
(b) Scatterplot documenting evaluation accuracy over-performance on long questions.
Figure 5: Longer synthetic questions have a measurable impact on evaluation accuracy over-performance. (a) Some models are brief, and some models produce overly long questions. (b) Longer synthetic questions correlate with over-performance on the generated evaluation benchmark.

### v2

@@ 4.2 Question generation models
Table 2: Question generating language models
Google gemma-3-4b-it [36]
gemma-3-12b-it [36]
gemma-3-27b-it [36]
Meta Llama-3.2-3B-Instruct [37]
Llama-3.1-8B-Instruct [38]
Llama-3.3-70B-Instruct [38]
Llama-4-Maverick-17B-128E-Instruct [39]
OpenAI o4-mini [40]
gpt-4.1 [40]
gpt-4.1-mini [40]
Microsoft phi-4 [41]
Qwen Qwen2.5-7B-Instruct [42]
To explore the impact of model capability on successful generation of synthetic data benchmarks we tested a variety of language models during the synthetic benchmark generation phase (see Table 2). All open-weight models are available through HuggingFace [43] and the closed-weight models were accessed through paid APIs. Additional results are reported for an ensemble of the bolded models. During benchmark creation. the same model is used both for topic extraction and question generation.
4.3 Evaluating against human questions
The reformatted human annotated question/answer pairs are used to compare the grounded synthetic data benchmark performance against the human. Ideally, the synthetic benchmark from grounding documents would produce the same LM capability measurements as the human-curated benchmark for any given model under evaluation. As such, to validate our approach, various LMs of differing sizes are evaluated on both the human and synthetic benchmarks. Comparison of pure benchmark evaluation accuracy is performed using the Pearson accuracy correlation between human and synthetic benchmark accuracy numbers. Evaluating the comparative rank ordering of models uses the Spearman rank correlation of relative model performance on the given benchmark dataset. These two measures compare both absolute performance alignment and relative rank order preservation/stability.
4.4 Generated question quality characterization
Per-question quality features and benchmark dataset-level features are computed to determine the overall quality of the synthetic benchmark. Owing to the low level of human oversight and annotation time for creating the synthetic data benchmarking, LMs extract these somewhat imprecise features. This relies on the idea that with enough guidance LMs can be good alternatives to human evaluation [16, 17] and borrows from standard practice LLM-as-a-Judge grading of open-ended benchmark responses [18, 19]. Llama-4-Maverick-17B-128E-Instruct extracts the following information about each generated question:
@@ 4.4 Generated question quality characterization
    question token count: number of tokens in the question
@@
    question clarity: how clear and comprehensible is the question in isolation without the associated context
@@
    question groundedness: how grounded is the question in the context
@@
    answer correctness: how correct is the answer given the context
@@
In addition to per-question measurements, question diversity is measured over each benchmark as the average cosine similarity between all questions (using sentence-transformers/all-MiniLM-L6-v2 as the embedding model). The answer_correctness and answer_explanation_validity numbers flag which generating models are incapable of producing diagnostically useful questions.
@@ 5 Results
All experiments and development were conducted on a single 8×H100 on-premise machine. Development consumed approximately twice the compute of final evaluations. Final compute was split between synthetic dataset generation and benchmark evaluation via Inspect-AI. The most expensive model, gpt-4.1, was only used during final results, processing 3.6M input and 1.1M output tokens ( \$16), with compute evenly divided between dataset creation and evaluation.
5.1 Answer correctness
Synthetic benchmark answers must be accurate, a task made easier via the grounding documents that supply the necessary information. To verify generated answer accuracy, an LM scored each question-answer pair and its justification (1-10 scale) using the full context. As shown in Figure 3, reasonably sized models scored above 9.5 in accuracy, with smaller models performing noticeably worse. While model size/capability generally correlates with desirable properties for synthetic data generation, notable exceptions (like gpt-4.1) are discussed later.
@@ 5.1 Answer correctness
Figure 3: Synthetic data generated question correctness and explanation validity on a scale of 1 to 10.
5.2 Question clarity and groundedness
@@
Figure 4: Generated question clarity and groundedness compared against the original questions.
Question clarity is inherently tied to difficulty, as simple questions have the potential to be clearer, easier to understand, and easier to answer. Conversely, a needlessly opaque question is harder to answer while providing minimal additional diagnostic relevance, though a model’s ability to reason through more opaque questions indicates general capability. This study found that most generated questions have acceptable clarity without significant variation between the generating models. Most generated synthetic questions are longer and more complicated than the original questions. This increased average question complexity combined with similar levels of question clarity suggests most synthetic questions are well written. Manual random spot inspection supports this as we observed few failures.
How grounded in the source material (context chunk) a given question is indicates how much outside information both the generation and evaluation models must leverage to correctly interpret what is being asked and answer. Low grounding indicates the question/answer pair cannot be found within the context chunk, or that significant external information is required. High grounding means all necessary information is found within the context chunk. Therefore the generated question is testing for information aligned with and present in the context. Figure 4 highlights the question clarity and groundedness scores for the question generating models. While most LMs generate similar clarity values, the trends in groundedness score are more complex. Some capable models (like Llama-4-Maverick-17B-128E-Instruct) score well, other capable models (like o4-mini) score worse, with smaller models in between (gemma-3-4b-it). We observed two grounding failure modes: either the model hallucinates irrelevant content absent from the context chunk (phi-4 commonly did this), or the generating model uses knowledge in its weights to expand upon the information presented in the context (o4-mini). See Appendix F for an example from phi-4. Models of medium size perform well on groundedness, simply reformatting information presented in the context.
5.3 Impact of question length
[…]
@@ 5.3 Impact of question length
(a) Question token count by generating model. Most models are within $20\%$ of human, but some produce overly long questions.
@@
(b) Scatterplot documenting evaluation accuracy over-performance on long questions. Models like gpt-4.1-mini plot well above the trend line.
Figure 5: Impact of question length on evaluation accuracy over-performance.
5.4 Benchmark question diversity
During synthetic question generation, the explicit topic extraction step provides a backstop against degenerate questions which only focus on the most obvious elements of a context chunk. To compare the variety and coverage of the generated questions we compute the average cosine similarity between all question pairs (excluding self pair) over the generated synthetic dataset. While most generating models produce similar diversity scores closely aligned with the human questions, some highly capable models (such as gpt-4.1) show comparatively worse question diversity while performing well in clarity, correctness, and groundedness. Figure 8 in Appendix C shows the question diversity per generating model along with the original question diversity.
5.5 Benchmark model accuracy correlation
Figure 6: Benchmark evaluation accuracy correlation generated by model ensemble demonstrating reasonable correlation between human and synthetic evaluation. Pearson accuracy correlation $0.745$.
Ideally, we want synthetic benchmarks to correlate well with the evaluation accuracy measured by human curated questions. This ends up being true for some generating models, but not all. Figure 6 shows the evaluation accuracy scatterplot comparing the reformatted human questions against the synthetic generated benchmark questions for the five bolded ensemble models in Table 2. This ensemble produces a Pearson accuracy correlation of $0.745$ between the evaluation accuracies for synthetic and human benchmarks. Each plot marker is the average evaluation accuracy for a given model-dataset combination. While this model ensemble produces reasonable overall accuracy correlation, occasionally specific LMs generate undesirable benchmarks. Figure 12 in Appendix E showcases a failure where the synthetic benchmark generated by gpt-4.1-mini contains overly easy questions where all models perform near $100\%$ evaluation accuracy (Pearson accuracy correlation $0.402$).
5.6 Benchmark model rank correlation
While direct Pearson correlation between human and synthetic benchmark evaluation accuracy is desirable, evaluations are often used to compare models. Therefore, a more important metric is the model rank order stability measured by Spearman rank correlation. Figure 2 plots the evaluation model performance across all datasets, comparing the synthetic and human accuracy Spearman ranking correlation of the evaluation model capability. Pearson accuracy correlation translates into Spearman ranking correlation and relative model order stability. Similarly, the overly easy gpt-4.1-mini synthetic benchmark in Figure 12 produces a mis-matched rank ordering (Spearman correlation $0.564$) in Figure 13(b). While ensembling various question generating LMs together will partially mitigate this effect, it is worth validating the set of generating models against the datasets in this paper to confirm the ensemble is not negatively impacting overall synthetic benchmark quality. The tendency of certain generating models to produce overly long, detailed questions that LMs find easier to answer may be the cause of the observed benchmark over-performance.
5.7 Multiple-choice vs open-ended questions
During this study, both multiple choice and open-ended questions were evaluated. None of our findings change whether using multiple choice questions for simplified evaluation benchmark grading, or open-ended questions to produce harder evaluation benchmarks where knowledge synthesis is required. The models that are good at creating synthetic multiple choice questions are also good at creating open ended questions. The models which create overly easy questions by including too much detail (like phi-4) do so for both open ended and MCQ, with the long question problem becoming even more pronounced for open ended questions as shown in Figure 9 in Appendix D.
Table 5 presents Pearson benchmark evaluation accuracy correlation and Spearman rank order correlation between the human and synthetic evaluations. See Appendix E for an expanded table showing correlations across all three LMs used for human question reformatting. Llama-4-Maverick was unexpectedly dominant in these final Table 5 results, even outperforming the ensemble of models (though not significantly).
Table 3: Pearson accuracy and Spearman rank correlation between synthetic and human benchmarks
Metric Pearson Accuracy Correlation Spearman Rank Correlation
Generating Model Multiple Choice Open Ended Multiple Choice Open Ended
Ensemble 0.7452 0.7095 0.9679 0.9209
Llama-3.2-3B-Instruct 0.4857 0.4709 0.6143 0.7143
Llama-3.3-70B-Instruct 0.6615 0.6034 0.9571 0.8813
Llama-3.1-8B-Instruct 0.6608 0.6217 0.7964 0.7275
Llama-4-Maverick 0.7891 0.7233 0.9679 0.9648
gemma-3-12b-it 0.6879 0.4085 0.8929 0.8198
gemma-3-27b-it 0.7477 0.6962 0.9571 0.9516
gemma-3-4b-it 0.6575 0.5787 0.6393 0.9209
gpt-4.1 0.6188 0.5216 0.9179 0.8549
gpt-4.1-mini 0.4015 0.5780 0.5643 0.8418
o4-mini 0.6542 0.5700 0.9250 0.7934
phi-4 0.1547 0.4965 0.6964 0.1780
Qwen2.5-7B-Instruct 0.5813 0.5905 0.7321 0.8901
5.8 Case study on two academic papers
To demonstrate our approach in a more realistic setting, we selected two academic papers: “Current Solutions and Future Trends for Robotic Prosthetic Hands” [44] and “Recent Advances in Large Language Model Benchmarks against Data Contamination: From Static to Dynamic Evaluation” [45]. The documents were chunked into 27 and 28 contexts respectively. “Current Solutions” averaged $7.7$ topics per context for a total of 210 questions. “Recent Advances” averaged $6.9$ topics per context for a total of 194 questions. For MCQ, the two synthetic benchmarks’ Spearman rank correlation is $0.911$, for OE the rank correlation is $0.969$. Under MCQ the gemma-3 models perform middle of the pack on these two benchmarks, but under OE questions the 12b and 27b outperform all other models. See Figure 14 in Appendix G for full rank ordering and examples of the generated questions.
5.9 Limitations
This work relies on a simple and possibly non-optimal document chunking approach. Extensions which cluster related topics or provide more fine-grained control over the information present in each chunk could improve the granularity and specificity of the generated benchmark. Including information found in tables, charts, and images would provide richer information sources that benchmark creation may take advantage of. Additionally, improving the nuance and difficulty of the questions significantly improves LM capability measurement. This likely requires careful management of the context information provided to the generating LM. Finally, the datasets used in this study to validate correlation with human written questions are all NLP extractive question answering type datasets. This limits the quality of the synthetic benchmarks as the grounding documents are fairly simplistic, especially compared to the two academic papers used in Section 5.8. Only the sec_qa dataset, which is grounded in an open source computer security textbook, comes close to the topic diversity and complexity of the example academic papers. All other validation datasets rely upon more general trivia type knowledge.
6 Conclusion
We introduce a methodology for creating synthetic benchmarks to evaluate language model (LM) capabilities grounded in high-quality document corpora selected by practitioners. This ensures the generated benchmarks are directly relevant to their specific domains and interests. This approach extends general-purpose human curated LM benchmarking into specialized domains, allowing for the identification of strengths and weaknesses that may not be apparent in broader evaluations.
Our method supports automatic generation of both multiple-choice questions (optimizing for ease of grading) and open-ended questions (optimizing for diagnostic relevance), allowing users to choose based on their evaluation goals. Validation against six human-curated NLP dataset achieves a Spearman rank correlation of $0.96$ and a benchmark Pearson accuracy correlation of $0.79$ between synthetic and human curated questions, suggesting that our approach is consistent with human-generated question/answer benchmarks. Additionally, we identify a potential synthetic data generation failure mode: very long questions can artificially inflate model performance, likely by overspecifying information related to the answer. This methodology enables the creation of targeted, relevant benchmark evaluations of domain-specific Language Model capabilities, facilitating deeper exploration into the capabilities and limitations of Language Models deployed for complex knowledge work.
Appendix A Expanded related works
This approach draws inspiration from NP-completeness, where easy to verify solutions may be hard to generate. The LM under evaluation has a harder job than the LM that converts existing documents into appropriately formatted test questions with accompanying answers. The evaluation LM must produce the correct information from just the question; by contrast, the benchmark creation LM has the source material. In other words, given the grounding paragraph, a language model should be much more capable at converting that textual information into a multiple choice or short answer question than it would be at zero-shot answering the very same question. This task difficulty discrepancy provides hope that this approach will produce viable results long term. Asking an LM to generate diagnostically useful evaluation benchmarks drawn only from information stored in model weights is difficult, but conversion of information from a high quality source into test questions is attainable.
This work draws on the rich history of language model evaluation, where LMs are already being used to judge and score the results of human curated benchmarks. When scoring outputs of an LM producing a multiple choice answer, it is often useful to have another LM reason about and extract the specified [A,B,C,D] value, instead of hoping the model under evaluation only produced the exact match answer letter [3, 4, 46]. This process becomes even more important when evaluating short answers from LMs instead of multiple choice responses, or when scoring the responses and providing reasoning as to why an answer was correct or incorrect [5, 6, 7]. This idea of automating question answer generation has been approached before. The most similar contemporary work is YourBench by Shashidhar et al. [8], which leverages the convergent idea of grounding documents supporting LM benchmark generation, demonstrating the synthetic replication of MMLU. Additionally, AutoBencher by Li et al. explores leveraging synthetic data (supported by privileged information) for model evaluation with RL to optimize the generated evaluation for discovering gaps in LM capability [9]. The use of grounding documents or privileged information greatly improves synthetic data benchmark utility, as the evaluation methodology no longer relies on simply asking an LM for questions (and answers).
Question from context and answer
Early Question Answer Generation (QAG) approaches attempted to solve the question creation from a context + correct answer input. These approaches include improvements in selecting better LM samples [47], increasing sample diversity over distributional shifts [48], explorations of interrogative prompting against the context [49], and evaluating round trip information consistency [50]. These early methods sometimes fine-tune generation models to improve task performance, though largely that is only done with smaller models [51, 52]. Some work focused on fine-tuning specialty models to improve option generation for multiple-choice questions [52]. Other early work during the shift to general purpose LMs involved leveraging those language models directly as data generation sources. This took the form of prompting the model to write questions (without grounding context) [53, 54, 55, 56], or configuring an LM-as-evaluator/interrogator dynamically interacting with the LM under evaluation to explore capability in a reference-free manner [53]. These approaches attempt to address the dataset leakage problem where benchmarks, once published, become part of the training dataset for the next round of models. As the capability of LMs increases over time, setting up adversarial LM-evaluator vs LM-evaluatee systems is becoming more feasible [57]. Nevertheless, extending those systems such that the LM-evaluator has grounding documents to pull from should strengthen the performance.
Scaffolding question generation with grounding documents
With the rise of general purpose token-tumblers like ChatGPT, longer and more complex grounding documents become increasingly useful. There is no longer a need to constrain synthetic benchmark creation to a short paragraph context + correct answer input; the LM can operate on larger and more complex document chunks, with less pre-processing required to make the input context data useful. Contemporaneous work YourBench [8] and AutoBencher [9] can leverage lightly-filtered web-pages as the context, chunking the data into rough paragraphs for ingestion into the question generation and grounding system. The SecQA benchmark was created partly through synthetic data grounded in an open-source computer security textbook [10]. Similarly (though much earlier), the PubMedQA benchmark was created through human and synthetic data generation grounded in biomedical research paper abstracts [11]. Some research demonstrates that operating purely on document summaries produces inferior generated questions compared to leveraging the full document [12]. With modern long-context LMs, summarization is becoming less of a problem, as models can manipulate significant volumes of text within their attention context. Synthetic data question generation systems have seen rapid improvement. This is partly driven by capability improvements in the LMs powering synthetic data generation. However, there has also been exploration into improvements in prompting for diverse question generation [13] and breadth of topic generation for chatbot arena style questions [14]. Enhancing retrieval augmented generation (RAG) systems for domain specific applications has also used synthetic data question generation to compare baseline RAG to fine-tuning [15].
LM as feature extractors
There are many aspects of text that one might want to quantify, however imprecisely. However, devoting the required copious human annotation hours is just impractical. Current LMs end up being acceptable feature extractors, when provided detailed prompts. Chiang and Lee demonstrate that LMs can substitute for human evaluators in assessing text quality with strong correlation in the results [16]. Similarly, Han et al. leverage LM evaluators for judging quality and find that results correlate highly with human judgments [17]. The use of LMs to extract features from text is not to dissimilar from using the LM-as-a-Judge [18] pattern for large scale open ended capability evaluations [18, 19, 21]. The LM is capable of grading open ended evaluation benchmark questions for correctness; therefore determining features other than answer correctness from a text sample is viable. However, there are limits on what LM-as-a-Judge is capable of ground truthing, especially in fields like information retrieval [20, 21] where LMs under-perform on more complex relevance judgments. Relying on LMs to extract properties like answer correctness given sufficient scaffolding and context is a well-used pattern.
Multiple choice vs open ended
When constructing knowledge or fact based evaluation benchmarks, there are two broad categories of answers: multiple choice and open ended. Open-ended questions provide the model less information and require synthesis [22], but at the cost of needing another LM to evaluate the correctness of the results (LM-as-a-Judge [18][23]. Multiple choice question (MCQ) results can be parsed and graded more simply. Modern evaluation frameworks like Inspect-AI [26] enable shuffling answer options and contain default prompts for both open-ended and MCQ evaluations. With MCQ, having the answers visible to the LM can artificially inflate performance [24], even going as far as the model identifying the correct answers without seeing the question [25]. Both open-ended and multiple choice questions have weaknesses to question reformatting and phrasing where semantically identical questions produce different answers [27]. The simplest form of open-ended evaluation relies on a judge LM, though other frameworks have been proposed which attempt to improve the precision of open-ended evaluations [28, 29].
Appendix B Impact of question and answer length